OpenStudy (aravindg):

let f be a real fn satisfying f(x+y+z)=f(x)f(y)f(z) for all real x ,y ,zif f(2)=4 and f'(0)=3 find f(0) and f'(2). @Physics

OpenStudy (aravindg):

james help

OpenStudy (jamesj):

is that your pic btw?

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

hw z it?

OpenStudy (jamesj):

smoking

OpenStudy (aravindg):

lol help me in this

OpenStudy (aravindg):

.........................james?

OpenStudy (jamesj):

Ok. Thinking about this a bit.

OpenStudy (aravindg):

.........

OpenStudy (jamesj):

Well, we need to evaluate f and f' and 0 and 2 obviously. What's not obvious to me right now is how to relate these quantities at x = 2. But for x = 0, it's not so bad. For example, f(3x) = f(x)^3 and hence 3 f'(3x) = 3 f'(x) f(x)^2 or f'(3x) = f'(x) f(x)^2 Hence f'(0) = f'(0) f(0)^2 => f(0)^2 = 1 => f(0) = 1 or -1

OpenStudy (aravindg):

othr one?

OpenStudy (aravindg):

.,..

OpenStudy (jamesj):

Let z = 0 and x = y = 1. Then f(2) = f(1)^2 . f(0) Also, if we let x = y, with z = 0, then f(2x) = f(x)^2 . f(0) hence 2 f'(2x) = 2 f(0) . f'(x) . f(x) ------ (*) Again, not yet clear to me yet how to pull f'(2) from these.

OpenStudy (aravindg):

hey i didnt get how u found f(0)

OpenStudy (jamesj):

You agree that: f'(3x) = f'(x) f(x)^2 ?

OpenStudy (aravindg):

how?

OpenStudy (jamesj):

By differentiating both sides of f(3x) = f(x)^3

OpenStudy (aravindg):

:( i didnt get whhere is y and z

OpenStudy (jamesj):

Alternatively with that equation: f(3x) = f(x)^3 set x = 0 then f(0) = f(0)^3 => f(0) = -1, 0, 1 ** I set x = y = z hence f(x + y + z) = f(3x) and f(x)f(y)f(z) = f(x)^3

OpenStudy (aravindg):

can i do that?

OpenStudy (jamesj):

Sure; why not. x, y and z are arbitrary. I have to go and have breakfast with friends. I'll think about the other part of this problem and post what I get.

OpenStudy (aravindg):

when???

OpenStudy (jamesj):

As soon as I have the answer. It bugs me when I don't have the answer, so I won't be able to fully stop thinking about it.

OpenStudy (aravindg):

wel can u be online within 15 minutes?

OpenStudy (jamesj):

I'll try, but I don't promise.

OpenStudy (aravindg):

:( plzzzzz

OpenStudy (sriram):

Let z = 0 and x = y = 1. Then f(2) = f(1)^2 . f(0) we got f(0)=+1 or -1 and we know f(2)=4 hence 4=f(1)^2 *1 or 4=f(1)*-1 as f(1)^2 > 0 hence f(0)=+1

OpenStudy (aravindg):

helpppppppppp

OpenStudy (jamesj):

@sriram: that makes sense.

OpenStudy (aravindg):

2nd part

OpenStudy (jamesj):

Can you confirm that "f(2)=4 and f'(0)=3" is what the problem says?

OpenStudy (aravindg):

no

OpenStudy (jamesj):

So if it doesn't say that f(2) = 4 and f'(0) = 3, what does it say?

OpenStudy (jamesj):

The reason I ask is that you'd think this function f is something like \[ f(x) = e^{ax} \] But with those values, I can't make that work. However even setting that form of f aside, I'm running into other difficulties trying to solve this with those values.